Number of paired electrons in O2 molecule is :
a. 7
b. 8
c. 16
d. 14

Grade:upto college level

1 Answers

Deepak Patra
9 years ago

(d) O2 – Oxygen (Z = 8) has following molecular orbital configuration of O2.

O2 (16e-) –s 1s2, s* 1s2, s 2s2, s* 2s2, s2px2, {p 2py2 = p 2pz2, {p* 2 py1 = p* 2pz1 i.e., 2 unpaired and 14 paired electrons.